OpenStudy (anonymous):

[2.06] Solve the equation Quantity x minus 6 all over 2 = 3 for x. 0 12 11 6

OpenStudy (anonymous):

Okay so you have (x-a)/b=c right? First step I would suggest is taking the b up to the otherside and multiplying so you get: x-a=b*c okay? Then to solve for x, you just move the a to the otherside So ultimately x=(b*c)+a I'm trying not to give you the answers,but instead help you understand it as a general formula, I hope that is okay?

OpenStudy (anonymous):

dont really get it

OpenStudy (anonymous):

Okay no problem So if you have (x-6)/2=3 you can times it all by 2 so you get x-6=3*2 so then you know that x-6=6 okay? so if you want 'x' you would rearrange it so that x=6+6? Does that clear things up?

OpenStudy (kainui):

supanick, maybe try guessing at how to turn the words into an equation and we can help you figure out if it's correct and why.

OpenStudy (anonymous):

yes it did does thanks